Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News

  • First 28 antenna science observation (30 Jul)
  • Double maintenance day Tuesday-Wednesday.
  • No general commissioning meeting this Thursday; there will be a small group meeting under Vivek to discuss Reference Pointing status and testing plans.
  • Next Thursday Jim Jackson will do a walk through of the antenna electronics block diagram.
  • Key dates:
    • 15 Sep: Call for Proposals
    • 01 Oct: Proposal deadline (A config; lowband for end of B config)
  • Staff Travel:
    • Dickman: 02 Aug - 04 Aug
    • Perley: 03 Aug - 09 Aug
    • Hayward: 10 Aug - 30 Aug
    • Langley: 13 Aug - 15 Aug
    • DRAO: Carlson: 16 Jul - 19 Aug
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Bob Sault: 26 July - 27 August (?)
    • Brent Carlson: 23 August - 04 September (tentative)
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)

Critical Commissioning review/reports

  • Overall status: Working to complete RSRO/ECSO projects

  • Highest priority development areas; readjusted based on resolution of several outstanding OSRO blockers.:
    • 1) RSRO/ECSO low frequency science
    • 2) OSRO/RSRO blockers (focused on missing data/data quality issues)
    • 3) RSRO/ECSO high frequency science as possible
    • 4) CSV testing (switched power, etc)
      • Wednesday will be dedicated to 3-bit sampler testing to advance progress.
    • 5) Demo science observing
    • 6) OSRO science observing

  • Note: scheduling flexibility adapted for remainder of D config.

System Verification/Hardware Commissioning:

Testing and characterization of the array based on the design goals.

Targets

Area Target Priority JIRA Deliver Status
High Frequency Reference pointing quantitative analysis 0 ECSV-16/ECDATA-91 2010Q3 Issues uncovered - see report below
Calibration Phase transfer 0 ECDATA-271 2010Q3 Needs better weather for real conclusions.
Calibration Switched power gain correction 0 - 2010Q3 pending SDM implementation for further work
Calibration RFI Sweep for all bands 2 - 2010Q3 Pending wide band widths
Calibration Develop set of frequency accuracy tests (X band is complete) 2 - 2010Q4 Perley:
Calibration Delay clunking 2 - 2010Q4 Rupen/Pokorny:

Reports

  • Antenna: Perley/Sowinski; note, the following are equipped with 3-bit samplers: EA15, EA22, EA27, EA28
    • EA20 encoder error (unresolved). Resolved.
    • EA02 spikes (unresolved).
    • EA09 IF: C/D (LCP)
    • EA10 L,X,Q not working.
    • EA11 IF: B issues; Kerry is planning to swap the deformatter.
    • EA16 IF: A/B (RCP)
  • Receiver: Perley
    • Ken on new receivers: first wideband X was installed on EA07
  • Array: Perley
    • Spurious flagging (MCAF) issue status: MR: 2 issues; 1st is MCAF problem where flags are not turned off (affects large sections/all of the array); fix is ready to be put in (Wednesday). 2nd is good data getting flagged; not as well understood; looking at issues with the subreflector; tests will be set up to move subr back and forth to examine repeatability and precision.
  • Correlator: Rupen
    • CBE issues relating to missing scans has been largely solved by the 30Jul version; now have sets of 'death' scripts to excite issues and test for solutions/recurrence. Unfortunately there are less predictable problems with the new version which require manual intervention; Martin continues to have the system through Tuesday to investigate; exploring memory management issues (no magic bullet but a systematic review/correction of the existing handling). Able to fall back to 02Jul version as needed for stable low frequency RSRO.

Observing Band Checkout

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

Targets

Reports

  • Low Frequency (L/S/C): Momjian
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-38

  • Mid Frequency (X/Ku): Dhawan
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links

  • High Frequency (K/Ka/Q): Sjouwerman/Claussen
    • Status:
      • Group meeting Thursday led by Vivek on reference pointing.
      • Operating; users should include one of the canonical flux calibrators in SB
      • Reference pointing in C or X
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-30

  • Update on Switched power

Observing Support:

Testing and development of the end-to-end capability of the array to support observations to the community (Documentation, OPT, Scheduling, Instrument models, Calibrator survey/models DB, Rapid response observing, demo science, post-processing).

Targets

Area Target Priority JIRA Deliver Status
Software CBE averaging/decimation (time/freq) 1 - 2010Q3; vci averaging - honors other-than-1s dumps.  
Software/Calibration Recommended data path for weights in calibration 1 - 2010Q3  
Flux Calibration SBs for Flux calibrator models 1 ECSV-43 2010Q3 Needed for D config (2 freqs/band); 5 h of observations ready

Reports

  • Meta summary
    • Multi-config SBs in regular use; note 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

  • Software Systems (Overlord: Butler):
    • Versioning Page: https://staff.nrao.edu/wiki/bin/view/EVLA/EVLASoftwareVersioning
    • OPT (Lorant): 1.05 release today. Nicely organized testing campaign somewhat executed late.
      • Planetary SB fail; reference pointing fail. BB: OPT is okay; model2script will be reworked; retain production version for now; exploring issue further today.
    • Scheduling (Joan): In use. Beginning to assign priorities to CSV testing/demo science.
    • Executor (Barry): Investigating phase jumps associated with small frequency changes (e.g., Doppler setting); no change in observing strategy (must calibrate before/after). JJ: Looking at the synthesizers for issues.
    • Corr (CBE/CM; Michael): Status of 30Jul version (as above). CM has had some slight changes to support the switched power implementation.
    • Archive (John): No updats.
    • Post-processing
      • CASA (Nick): cvel problems with AW775; high priority. Need report/update.
      • AIPS (Eric): Working on spectral line version of PCAL.
    • ASDM
      • SDM total power/switched power table write-up (Rupen).
      • Flag table write-up (Myers/McMullin).FBT in progress.
      • SDM document (with at least table descriptions/contents/units). Re-scheduled for this month; taking the opportunity to push forward some needed operational modifications.
      • Priority 3: Ephemeris table resolution (Butler).
      • Precise UVW calculations in filler (McMullin/Viallefond). Being reviewed.

  • Rapid Response Support (Frail): TRAP

  • Flux Calibrator Models (Amy/Rick): TCAL
    • Report: TCAL0001_sb1576498_1.55377.02314818287, TCAL0001.sb1575783.eb1620753.55378.450005439816. 5 hours of observations taken (low frequencies). Amy report - pending.; Some discussion of need for good models for polarization calibrators; needs the outstanding polarization calibrator list (Myers).
    • Polarization calibrator list (Myers/Perley)
    • Links: ECSV-43

Observing Mode Support:

Testing and development within broad observing modes distinguished by their science goals (continuum, spectral line, polarization).

Targets

Area Target Priority JIRA Deliver Status
Spectral Line Bandpass stability 1 ECSV-35 2010Q3  
Spectral Line Doppler tracking 1 ECSV-32 2010Q3  
Spectral Line Spectral line setups; calibration procedures recommendations (OSRO/RSRO) 1 ECSV-31 2010Q3  
Polarization On-axis leakage calibration 1 ECSV-25 2010Q3  
Polarization R-L phase calibration 1 ECSV-25 2010Q3  
Polarization Observer Recommendations (OSRO/RSRO) 2 ECSV-25 2010Q3  
Polarization Beam squint calibration 2 ECSV-25 2010Q3  
Polarization Primary beam (off-axis) calibration 2 ECSV-25 2010Q3  
Polarization Polarization calibration service observation plan 3 ECSV-25 2010Q4  
Demo Science SBs for approved EPO projects 1 ECSV-57 2010Q3 Started for a subset

Reports

  • Spectral Line Modes (Ott/Goss): TSPE
    • Testing:
      • Report on Doppler setting/tracking recommendations (Ott). Bug found in CASA cvel (see CASA later).
      • Spectral Line general commissioning tests (diff spw widths, diff frequencies, bandpasses, etc): ECSV-92
      • Bandpass stability: ECSV-35:
        • TSPE0005_sb1261338 (3C84) ready - but day time
        • TSPE0005_sb1661507 (2253+161) Ready - scheduled for this week.
    • Links: ECSV-19
  • Polarization Modes (Myers): TPOL
    • Status:
    • Testing
      • Report on off-axis polarization calibration (Heesen) - needs ticket to submit to.
      • Multi-band leakage/pos. angle calibration report pending (Myers)
      • Report from Josh on stability measurements making use of test data sets (X_osro et al); reports will be posted to JIRA shortly.
    • Links:

  • Demo Science (Chandler/Condon): TDEM
    • Status: Begun; list requires prioritization; limited remaining time in D-config; Will try for IRC+10216 this week as possible.
    • Testing
Project StatusSorted ascending Time Comments SB
IRC+10216 Blocked 33 K, Ka, Q spectral sweep; no recirculation; pending test Not submitted
M87 Blocked 1x6 Ka TDEM0006_sb1470572 (not submitted)
Crab nebula Ready 1x10 L-band OSRO1/OSRO2 TDEM0007_sb1524351
M51 Ready 1x8 L band; needs corr test TDEM0008_sb1542082
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Not submitted
M82 Ready 1x2 L band TDEM0010_sb1660855

  • Advanced: Phased Array (Lazio/Deller): TPHA
    • Status: Need an update to FPGAs to support phasing (previously removed). BB: Some planning for Ransom/Demorest to visit for some pulsar tests (noted Mark V's coming in mid-September).
    • Testing: quiescent
    • Links: https://staff.nrao.edu/wiki/bin/view/EVLA/PhasedArray

  • Advanced Planetary: TPLA
    • Testing
      • Cassini observations obtained.; StillUnder review by Bryan.
    • Status:
      • OPT handles planetary observations now (following fix) using internal ephemeris; still needs polynomial tracking in the executor; handling of near-field sources in the SDM (Butler/Viallefond).
  • Advanced: Solar: TSOL
    • Status: Planning/prototype testing ongoing; working with L, Ku (special Rx setups in these two bands). No progress. Awaiting Rick/Bob returns. Noted that there is a CME scheduled for tonight; between maintenance day so no disruption to observing.
  • Advanced: Pulsar (Brisken): TPUL (See Phased array above for brief update.)
  • Advanced: VLBI (Walker/Brisken): TVLB
  • Advanced: Astrometry (Mioduszewski)

  • RSRO/ECSO: TRSR
    • Status: RSRO Readiness (Cycle1) Master List
      • Need RSRO/ECSO SBs in now! Please note the following:
        • Follow the nominal guide for OSRO (e.g., ensure wrap is set for first few scans, dummy scan for each frequency change).
        • Need 2x2 minute OSRO scans at the beginning of the SB to provide additional time for the correlator to configure (priming for RSRO); it's recommended that these be at a different frequency than the science observations to enable easily distinguishing them from the science data in post processing.
        • No scans longer than 90s (or so - exact number TBD).
        • Set it up using the extended OSRO configurations
    • RSRO/ECSO Science Observations (Red=recent observations requiring report; Blue=active; Black=pending):
Project Band(s) Test Code Test Status Allocation Status (D) Notes JIRA
AA30 Q TRSR0003 - 0/51 Blocked RSRO-1
AB1345 L TRSR0004 21May10 13/50 In Progress RSRO-2
AB1346 K TRSR0005 23Jun10 0/42 Blocked; 3 hrs failed RSRO-3
AC982 CKKaQ TRSR0002 28Jul10 9/64 Blocked (high freq) RSRO-4
AH1006 LCXKa TRSR0006 - 8/36 Blocked (high freq) RSRO-5
AK726 Ka TRSR0007 - 0/43 Blocked RSRO-6
AL746 CKa TRSR0008 05Jul10 1.5/11.4 Blocked (high freq) RSRO-7
AM1014 C TRSR0009 22Jun10 0/0 C config RSRO-9
AS1015 C TRSR0010 21May10 4/4 Complete
1482295,1626830
RSRO-11
AT374 C TRSR0011 25Jul10 0/0 C config RSRO-12
AV318 L TRSR0012 - 12/12 Complete; PI converted to OSRO
1491201
RSRO-13
           
AW775 L TRSR0017 05Jul10 16/72 In Progress
1660006
-
AM1005 KKaQ TRSR0014 - 0/20 Blocked -
AM1038 Q TRSR0019 21May10 0/10 Blocked -
AC992 L TRSR0018 26Jun10 4/4 Complete
1426692,1427000
1427915,1705655
-
AO255 Ka TRSR0016 - 0/70 Ready
1423157,1466135
1467049,1467331
1481381
-
AM1018 Ka TRSR0015 - 0/48 Blocked -
AB1353 Ka TRSR0013 - 0/51 Blocked -

    • Resident Schedule: Posted on home page of twiki now!
      • 01 Jan - 16 May: Adam Deller (AC982: RSRO-4): 300A
      • 31 Jan - 31 Jul: Laura Chomiuk (AS1015/AK726: RSRO-11/RSRO-6): 200C
      • 04 Feb - 01 Jun: Jacqueline van Gorkom (AV318: RSRO-13): 200D
      • 25 Feb - 01 Jun: James Miller-Jones (AM1014: RSRO-9): 200E
      • 02 Mar - 07 Jun: Manuel Aravena (AA330: RSRO-1): 300G
      • 16 Mar - 08 Apr: Dave Green (AB1345: RSRO-2): 300G
      • 20 Apr - 30 Jun: Crystal Brogan (AB1346: RSRO-3): 200B
      • 20 Apr - 30 Jun: Todd Hunter (AB1346: RSRO-3): 200B
      • 30 Apr - 15 Jul: Joseph Lazio (AC982: RSRO-4): 266
      • 07 Jun - 16 Sep: Laura Perez (AC982: RSRO-4): 200E
      • 28 Jun - 05 Oct: Volker Heesen (AH1006: RSRO-5): 200G
      • 05 Jul - 26 Oct: Russ Taylor (AT374: RSRO-12): 301; Shannon Jaeger is also here.
      • 21 Jul - 23 Sep: Dave Green (AB1345: RSRO-2): 300G
      • 15 Sep - 15 Dec: Adam Leroy (AL746: RSRO-7): 200C
      • 01 Oct - 01 Jan: Hendrik Linz (AC982: RSRO-4): 200C
      • 01 Jan - 01 Apr: Jane Greaves (AC982: RSRO-4): 300G
      • 02 Jan - 02 Apr: Stuart Corder (AC982: RSRO-4): 300G
    • Current Residents
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
        • This week:
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
        • This week:
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
        • This week:
      • Peter Hofner (POC: ); area(s):
        • This week:
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
        • This week:
    • Past (and future residents)
      • Joe Lazio (POC: Claire Chandler); area(s):
      • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
      • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
      • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
      • Manuel Aravena (POC: Chris Carilli); area(s): high frequency observing
      • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
      • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
      • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)

  • Science Support Team: Adding assignments for targeted data reviews
    • McMullin: testing (regressions), operations (nightly scheduling), CSV organization, ASDM issues, PEPs
    • Myers: vacation; SKA
    • Ott: travel
    • Greisen: PCAL
    • Dhawan: Reference pointing
    • Kogan: UVRFI testing with Frazer
    • Rupen: CBE work with Martin; 3-bit sampler testing, all things correlator
    • Krauss: flagdata interface conclusion; web page documentation; pipeline regression script

Advanced Science Support:

These are the testing/development areas needed to fulfill the full capabilities of the EVLA along with additional observing modes for other branches of the community.

  • Lowband (Owen)
    • Status: Update to OSS for next proposal cycle; current plan is to put 6 dipoles on the non-moving antennas for the D->DnC configuration (2 per arm); put on Monday, take down on Friday; goal is to have a better idea of information to provide to users; BB: May have more impact than anticipated - thought we could hand edit scripts; might be more complicated - action to investigate with Bryan.
    • Receiver/feed tests status:
    • Calibration Procedures
    • Links: https://staff.nrao.edu/wiki/bin/view/EVLA/Lowband

  • Algorithms (Owen/Bhatnagar)

  • Astrometry (Mioduszewski)
    • Status: Wait for A array(?)

  • HPC
    • Status: Minutes from last meeting. FO: Noted most of the initial equipment is in hand; need commissioning plan, i.e., planning for the cluster to come; BB noted that the remaining cluster nodes are not likely to be purchased before 1 Oct (change from recent reports). RT: How portable will the system be (i.e., could it be used on cluster in Calgary?) - BB: Yes, that's the idea - very portable.
The Postprocessing Hardware group met on July 23 at 1pm.

Attending in Socorro: Chandler, Elias, Greisen, McMullin, Owen, Robnett; in Charlottesville:  Cotton, Halstead, Hatz, Hunt

  James reported that he has on order all the components for the EVLA cluster, except for most of the nodes. He has ordered two nodes for tests over the next few weeks before deciding on the details of final purchase of the remaining 15 nodes. This process should be complete in August. The focus of EVLA activities now needs to be in two areas: 1) the parallelization and optimization of the IO efficiency of the CASA code  and 2) the details of the procedures for setting up and queuing jobs for the  cluster, i.e. the commissioning of the cluster.

  Nick reports that more work is needed to develop a plan for parallelization of the CASA code. James suggests that a technical meeting, possibly as often as every week, is needed for progress in the general HPC area. James will organize such a meeting.

   A plan is also needed for commissioning the EVLA cluster, supposedly involving Gustaaf and Joe.

  The group would like to continue to have this Charlottesville/Socorro postprocessing Hardware coordination meeting  but on its former once-a-month schedule.  We will schedule the next meeting as usual on the third Friday in August (August 20th at 1pm MDT).
    • Links:

AOB

Misc Actions

  • [McMullin/Frail]: Write up disposition of RR items.
  • [McMullin/Butler/Rupen/Owen]: Resolve discussion of requirements for DnC test of 4-band.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 02 Aug 2010
Topic revision: r4 - 2010-08-04, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback